OpenStudy (anonymous):

Find the point slope form of an equation of the line through the two ordered pairs (4,28) (16,98)

OpenStudy (anonymous):

find the slope first: m=(y-y1)/(x-x1)

OpenStudy (amistre64):

subtract points to determine slope (16,98) -(4,28) -------- 12, 70 ; slope = y/x = 70/12 = 35/6

OpenStudy (amistre64):

y-28 = (35/6)(x-4)